OpenStudy (anonymous):

Determine whether each statement is true or false. Explain. The number of permutations of two items from a data set is always two times the number of combinations when taking two objects at a time from the same data set.

jimthompson5910 (jim_thompson5910):

Since C(n,r) = ( P(n,r) )/(r!) and 2! = 2*1 = 2, this statement is true.

OpenStudy (anonymous):

oh ok thanks jim_thompson5910